# Remote Rowhammer Attack using Adversarial Observations on Federated Learning Clients

### v1

@@ III-B Evaluation Metrics
III-B1 Notations
To evaluate and validate the proposed attack strategy, we map the repeated rate of sparse client updates to the number of Rowhammer operations required for DRAM. We define the following terms:
  • $H_{\text{max}}(p)$: The maximum number of sparse updates processed within a single DRAM refresh period for an update fraction $p$ that leads to update size $S_{\text{update}}(p)$.
  • $p$: The fraction of model parameters updated in each sparse update.
  • $T_{\text{flip}}(r)$: The Rowhammer threshold, measured as the number of updates required to induce a bit-flip in row $r$.
  • $\overline{T_{\text{flip}}}$: The average Rowhammer threshold across all rows.
@@ III-B2 RL Performance Metrics
For the RL agent, we define the following metrics that comprehensively quantify the clustering behavior and consistency of induced parameter updates:
  • Repeated Update Rate (RUR): Measures consistency of parameter indices updated across episodes. Higher RUR indicates persistent, targeted updates:
    $\mathrm{RUR}=\frac{\sum_{t=1}^{T-1}|U_{t}\cap U_{t+1}|}{\sum_{t=1}^{T-1}|U_{t}|},$ (1)
    where $U_{t}$ is the set of parameter indices updated at episode t.
  • Cluster Density (CD): Evaluates how tightly parameter updates cluster within the model’s global parameter index space. Lower CD signifies highly localized updates conducive to Rowhammer effects.
    $\mathrm{CD}=\frac{L}{|\bm{\theta}|},\quad\text{where }L=\min_{1\leq j\leq k-m+1}(i_{(j+m-1)}-i_{(j)}+1)$
    denotes the smallest span containing 90% of updates in each episode.
This metric captures how frequently updates concentrate on a single layer. A higher value indicates that one layer receives a disproportionate number of updates.
III-B3 Bit-Flip Validation
Given the DRAM bandwidth $BW$, the refresh period $\Delta t_{\text{refresh}}$, and a fraction $p$ that determines the size $S_{\text{update}}(p)$, the maximum number of updates is computed as
$\ H_{\text{max}}(p)=\frac{BW\times\Delta t_{\text{refresh}}}{S_{\text{update}}(p)},$ (2)
where $S_{\text{update}}(p)$ is the size of a sparse update when a fraction $p$ of parameters is transmitted.
To assess the practical feasibility of inducing Rowhammer conditions, we estimate the number of expected activations on a single targeted DRAM row (or closely spaced group of rows) based on empirically measured update clustering behaviors.
The expected number of row activations induced by the repeated updates, denoted as $\mathbb{E}_{Activation}$, can be approximated as:
$\mathbb{E}_{Activation}=\text{RUR}\times H_{\text{max}}(p),$ (3)
where RUR is the empirically measured repeated update rate
Finally, given a sparse update fraction $p$, bit-flips are considered successfully triggered when $\mathbb{E}_{Activation}$ is higher than $\overline{T_{\text{flip}}}$.
@@ III-C Target System
The target system is a self-supervised federated voice recognition system, to emulate large scale voice assistant pre-training systems, which are widely deployed on various platforms including smart home devices, mobile phones, and other IoT devices, and involving large quantity of clients.
@@ III-C1 Model Structure
We train the agent with target FL ASR systems of the following five model structures, the size details of these models are presented in Table I.
@@ III-C2 Dataset
For dataset, the FL takes the Common Voice 17.0 dataset [39] as the training data, which is a large-scale multilingual voice dataset. We only use the English subset for simplicity. The subset contains 1.1M voice clips with a total duration of more than 3k hours. The data is randomly split into equal client partitions, where the size of the partition is determined by the number of clients.
@@ III-C4 Clients Setups
On the client side, each client emulates different physical environments via applying random Gaussian noise in addition to the adversarial noise generated by the RL agent, after which the mixed waveforms are resampled to 16kHz to match the Common Voice dataset. Also, the clients are set to send sparse updates, i.e., at each communication round, only a controllable portion of the most significant gradient updates are sent to the server.
@@ III-D RL Agent Design
We propose a two-stage PPO-based agent to generate adversarial waveforms that consistently induce parameter updates clustered within targeted indices of an ASR model’s parameter vector, denoted by $\bm{\theta}\in\mathbb{R}^{M}$. The two-stage design initially provides explicit guidance for parameter localization (Stage 1), followed by generalized adversarial capability under limited supervision (Stage 2).
@@ III-D1 Observation and Action Spaces
  • Observation: $\mathcal{O}_{1}=\{x_{\text{clean}},\mathbf{u}_{t}\}$, with binary vector $\mathbf{u}_{t}\in\{0,1\}^{M}$ indicating recently updated parameters.
  • Action: $\mathcal{A}_{1}=\{\delta\in\mathbb{R}^{D}\mid\|\delta\|_{\infty}\leq\epsilon\}$, constrained adversarial perturbations.
In Stage 2, explicit update signals are removed to prevent dependence on instantaneous signals, promoting robust generalization across diverse inputs:
  • Observation: $\mathcal{O}_{2}=\{x_{\text{clean}}\}$, relying solely on waveform input to generalize learned perturbation strategies.
  • Action: Maintained as $\mathcal{A}_{2}=\mathcal{A}_{1}$.
@@ III-D2 Reward Formulation
Target Specificity: Ratio term emphasizes updates within targeted parameter indices $[i:j]$.
Perceptibility Penalty:
$\mathcal{L}_{\text{perc}}=\lambda_{1}\|\text{STFT}(x^{\prime})-\text{STFT}(x)\|_{F}+\lambda_{2}\,\text{rms}(\delta)$ (5)
This perceptual constraint simultaneously limits frequency-domain artifacts and temporal waveform distortions [10].
TABLE II: Rowhammer Attack Thresholds under Different Patterns
Access Pattern Data Pattern (Victim) Data Pattern (Aggressor) Threshold (ACTs)
Single-sided 0x00000000 0xFFFFFFFF 240K
Single-sided 0xFFFFFFFF 0x00000000 185K
Single-sided 0x55555555 0x55555555 260K
Single-sided 0xAAAAAAAA 0xAAAAAAAA 265K
Double-sided 0x00000000 0xFFFFFFFF 140K
Double-sided 0xFFFFFFFF 0x00000000 115K
Double-sided 0x55555555 0x55555555 160K
Double-sided 0xAAAAAAAA 0xAAAAAAAA 165K
@@ III-D3 Policy Network Architecture
The policy network employs a hierarchical architecture with a shared feature extractor and stage-specific decoders. The network first processes raw waveforms through a multi-resolution convolutional frontend, followed by bidirectional temporal modeling, before diverging into stage-dependent perturbation generators.
Shared Encoder Design: The encoder processes raw waveforms through parallel convolutional branches with progressively larger kernels (3, 5, and 25 samples). These branches extract localized phoneme features (4kHz resolution) and broader spectral patterns (1kHz resolution), concatenated into a 1536-channel representation. A 4-layer bidirectional LSTM then models temporal dependencies across 500ms windows, matching common ASR systems.
Stage-Specific Generators:
  • Stage 1 explicitly incorporates parameter update locations by projecting the binary mask $\mathbf{u}_{t}$ into a 1024-dimensional embedding, concatenated with encoder outputs. This design follows curriculum learning principles, accelerating early-stage association between perturbations and targeted parameter updates.
  • Stage 2 replaces explicit masking with an 8-head cross-attention mechanism. A learned query vector $q\in\mathbb{R}^{512}$ interacts with encoder features through scaled dot-product attention, enabling the agent to focus on temporal regions most correlated with persistent parameter clustering. Both stages employ identical transposed convolutional blocks for perturbation synthesis.
III-D4 Training Protocol
Curriculum Phases: The phased training aligns with the policy network’s dual-head architecture to optimize parameter targeting:
  1. 1.
    Stage 1 (50k episodes): This phase trains the shared encoder and Stage 1 head to establish baseline correlations between waveform perturbations and localized parameter updates.
  2. 2.
    Transition (10k episodes): Observation masking $\mathbf{u}_{t}$ is progressively replaced with zero vectors. The cross-attention head is warm-started using Stage 1 encoder weights, with orthogonal regularization ($\beta=0.01$) applied to BiLSTM layers to prevent feature collapse.
  3. 3.
    Stage 2 (20k episodes): Full $\mathcal{O}_{2}$ observations with $\epsilon=0.1$ enhance attack potency. The attention query vector $q$ is fine-tuned while freezing the shared encoder’s first three BiLSTM layers, balancing generalization with temporal feature retention.
@@ IV-A Experimental Setups
[…]
@@ IV-B Rowhammer Test Bench
Considering that FL may use different memory replacement strategies and that data diversity plays a role, this work tested different Rowhammer attack thresholds, based on two access patterns and four data patterns. Compare to single-sided access pattern, the double-sided requires less activations. However, it is also more difficult to utilise as it requires two adjacent rows for victim data. The data pattern will be the data layout between victim rows and aggressor rows. The tests not only reveal patterns that are more likely to induce bit flips, such as 0 → 1 or 1 → 0, but also demonstrate relatively more challenging patterns, such as 0 → 0 or 1 → 1. Specially, we use hexadecimal strings to represent data patterns that fills the entire DRAM row. For example, 0xFFFFFFFF represents 32-bit data filled with all ones, while 0x00000000 represents all zeros. The obtained attack thresholds are derived from the average values across all banks of the tested DRAM module.
These activation thresholds ensure that the attack induces bit-flips in at least 95% of the DRAM rows. The inability to obtain thresholds for all rows is likely due to the presence of exceptionally resilient cells within the module, which results in extreme deviations in the average threshold. These thresholds are specifically presented in Table II. All threshold values fall within the theoretical limit of 1,309K activations and reserve enough dummy rows activations to bypass TRR. It can be observed that double-sided Rowhammer generally requires fewer activations than single-sided Rowhammer. In the worst case, 265K activations are sufficient to induce bit flips.
@@ IV-C PPO Agent Training
We empirically evaluate our proposed PPO-based adversarial framework to assess its effectiveness in consistently inducing clustered and repetitive parameter updates in federated Automatic Speech Recognition (ASR) systems under varying sparsity settings, which are suitable for Rowhammer exploitation.
IV-C1 Experimental Setup
@@
We selected five widely-used ASR architectures of varying complexity, parameter counts, and architectural patterns for evaluation, see in Table I the five models and their parameter statistics.
For each model, two levels of parameter-update sparsity were considered:
  • Medium Sparsity (0.1% updates): Represents scenarios where moderate communication efficiency is required without severely limiting model convergence speed.
@@
  • High Sparsity (0.05% updates): Reflects extreme communication efficiency conditions, simulating highly bandwidth-constrained or large-scale deployments.
@@
The PPO agent was trained to generate adversarial waveforms constrained by a maximum perturbation limit ($\|\delta\|_{\infty}\leq 0.1$), ensuring the adversarial audio samples remain stealthy and practically imperceptible. The training dataset employed was the English subset of Common Voice 17.0 [39]. Hyperparameters for the reward function were empirically optimized as $\alpha=1.0$, $\beta=0.8$, and $\gamma=0.6$, with perceptibility penalty weights set to $\lambda_{1}=\lambda_{2}=0.5$.
IV-C2 Baseline Methods
To provide clear insights into the effectiveness of our two-stage PPO approach, we compared it against two baselines:
  • Single-Stage PPO (PPO-SS): PPO trained skipping stage 1 for the same number of total episodes.
  • Unperturbed: FL operating without any external influences.
IV-C3 Evaluation Metrics
To evaluate our PPO agent, we measure their performance with the following metrics, previously defined in III-D.
  • Repeated Update Rate (RUR)
  • Cluster Density (CD)
  • Average Rowhammer threshold across all rows ($\overline{T_{\text{flip}}}$)
IV-C4 Results and Analysis
Tables III and IV present comprehensive evaluations of our PPO-TS method against baseline approaches across different model architectures and sparsity levels.
TABLE III: Repeated Tensor Overlap Rate (RUR, %) across all models and sparsity levels.
Model Method RUR (0.1%) RUR (0.05%)
Conformer-CTC-S PPO-TS 73.7 68.3
PPO-SS 65.1 53.9
Unperturbed 12.5 9.3
QuartzNet 5x5 PPO-TS 75.6 71.1
PPO-SS 64.8 57.2
Unperturbed 11.8 8.7
Citrinet 256 PPO-TS 71.9 66.4
PPO-SS 57.4 52.9
Unperturbed 13.9 10.2
Squeezeformer-XS PPO-TS 72.5 68.1
PPO-SS 61.4 54.4
Unperturbed 11.2 8.1
Eff. Conformer-CTC PPO-TS 71.1 66.1
PPO-SS 59.3 51.7
Unperturbed 13.2 9.6
Repeated Update Rate (RUR): As summarized in Table III, our PPO-TS approach consistently outperformed both the PPO-SS and Unperturbed baselines. At high sparsity (0.05%), PPO-TS achieved RUR ranging from $66.1\%$ (Eff. Conformer-CTC) to $71.1\%$ (QuartzNet 5x5), significantly surpassing PPO-SS (approximately $51\%$$57\%$). This substantial improvement underscores the two stage design to learn robust, long-term strategies for generating perturbations that repeatedly affect the same parameters.
TABLE IV: Cluster Density (CD, %) across all models and sparsity levels.
Model Method CD (%)
0.1% 0.05%
Conformer-CTC-S PPO-TS 1.2 0.6
PPO-SS 3.7 1.8
Unperturbed 19.2 15.4
QuartzNet 5x5 PPO-TS 1.5 0.8
PPO-SS 4.1 3.2
Unperturbed 18.6 14.8
Citrinet 256 PPO-TS 1.0 0.7
PPO-SS 3.4 1.9
Unperturbed 17.9 13.9
Squeezeformer-XS PPO-TS 1.6 0.9
PPO-SS 3.8 2.4
Unperturbed 19.5 15.8
Eff. Conformer-CTC PPO-TS 1.1 0.5
PPO-SS 3.6 2.3
Unperturbed 18.2 14.3
Cluster Density (CD): As detailed in Table IV, PPO-TS resulted in significantly lower CD values compared to baselines across all tested configurations. At the high sparsity level, PPO-TS achieved CD values as low as $1.0\%$$1.6\%$, marking approximately a 2–3× reduction compared to PPO-SS (around $3.4\%$$4.1\%$). Such extreme clustering effectively concentrates updates into tightly bound memory regions, dramatically increasing the probability of repeatedly activating the same DRAM rows.
These results demonstrate the efficacy of our approach in systematically inducing concentrated and stable update patterns necessary for facilitating targeted Rowhammer attacks in realistic FL settings.
IV-D Bit-Flip Validation

### v2

@@ III-B Evaluation Metrics
III-B1 Notation
To evaluate the proposed attack strategy, we relate repeated sparse parameter updates in federated learning to the number of effective DRAM row activations that may be induced within a refresh window. We use the following notation:
  • $p$: Fraction of model parameters included in each sparse update.
  • $S_{\text{update}}(p)$: Size in bytes of one sparse update when the update fraction is $p$, including both parameter values and indexing metadata.
  • $H_{\max}(p)$: Upper bound on the number of sparse updates that can be processed within one DRAM refresh period for update fraction $p$.
  • $\Delta t_{\text{refresh}}$: DRAM refresh period.
  • BWBW: Effective bandwidth available to sparse update processing.
  • $T_{\text{flip}}(r)$: Rowhammer threshold of row $r$, measured as the number of effective row activations required to induce a bit flip.
  • $\overline{T_{\text{flip}}}$: Average Rowhammer threshold across tested vulnerable rows.
@@ III-B2 RL Performance Metrics
As illustrated in Process-II in Fig 2, centred to the success of our attack vector is inducing consistent repeated and clustered sparse updates from the clients. For the RL agent, we use metrics that quantify the persistence and spatial concentration of the induced sparse updates.
  • Repeated Update Rate (RUR): RUR measures the overlap between updated parameter sets in consecutive episodes or rounds. A higher RUR indicates that the agent repeatedly targets similar parameter indices:
    $\mathrm{RUR}=\frac{\sum_{t=1}^{T-1}\lvert U_{t}\cap U_{t+1}\rvert}{\sum_{t=1}^{T-1}\lvert U_{t}\rvert}$, (1)
    where $U_{t}$ denotes the set of parameter indices updated at step $t$.
  • Cluster Density (CD): CD measures how tightly updated indices are concentrated in the global parameter index space. Lower CD indicates stronger spatial locality:
    $\mathrm{CD}=\frac{L}{\lvert\bm{\theta}\rvert}$, (2)
    where $\bm{\theta}$ is the full parameter vector, $k$ is the number of updated indices in one step, $m=\lceil 0.9k\rceil$, $i_{(1)}\leq i_{(2)}\leq\dots\leq i_{(k)}$ are the sorted updated indices, and
    $L=\min_{1\leq j\leq k-m+1}\left(i_{(j+m-1)}-i_{(j)}+1\right)$ (3)
    is the minimum span containing 90% of updated indices.
Together, RUR and CD characterize whether the agent repeatedly concentrates updates on a small and stable region of the model, which is a necessary precursor to repeated accesses to the same memory region.
III-B3 Bit Flip Validation
Given the effective bandwidth $BW$, refresh period $\Delta t_{\text{refresh}}$, and sparse update size $S_{\text{update}}(p)$, the maximum number of sparse updates that can be processed within one refresh window is upper bounded by
$H_{\max}(p)=\frac{BW\cdot\Delta t_{\text{refresh}}}{S_{\text{update}}(p)}.$ (4)
This quantity is an upper bound because it does not account for protocol overhead, software processing cost, cache effects, or other system bottlenecks.
To estimate the number of effective activations on a targeted DRAM row, we combine the maximum update rate with the measured repeated update behavior. Then the expected number of effective activations within one refresh window is approximated as
$\mathbb{E}[A]=\mathrm{RUR}\cdot H_{\max}(p).$ (5)
Finally, for a given update fraction $p$, the attack is considered capable of creating Rowhammer relevant conditions when
$\mathbb{E}[A]\geq T_{\text{flip}}(r)$ (6)
for a targeted vulnerable row $r$, or more conservatively when $\mathbb{E}[A]$ approaches empirically observed Rowhammer thresholds on the test platform. In practice, this comparison should be interpreted as a feasibility criterion rather than a deterministic guarantee of bit flips.
@@ III-C Target System
The target systems consist of two variant, a self-supervised federated voice recognition system, and a federated CIFAR10 classifier, emulating large scale IoT or voice assistant pre-training systems, which are widely deployed on various platforms including smart home devices, mobile phones, and other edge devices, contributing to a large quantity of clients.
@@ III-C1 Model Structure
Model structures of the target systems, as well as details including the corresponding sizes and precision of these models are presented in Table I.
@@ III-C2 Dataset
For dataset, the ASR systems take the Common Voice 17.0 dataset [40] as the training data, which is a large-scale multilingual voice dataset. We only use the English subset for simplicity, which contains 1.1M voice clips with a total duration of more than 3k hours. As for the computer vision target system uses the CIFAR10 dataset [41]. In both cases, the data is randomly split into equal client partitions, where the size of the partition is determined by the number of clients.
@@ III-C4 Clients Setups
On the client side, for the ASR systems, each client emulates different physical environments via applying random Gaussian noise in addition to the adversarial noise generated by the RL agent, after which the mixed waveforms are resampled to 16kHz to match the Common Voice dataset. For computer vision, the printing and re-digitalization of the dataset is emulated with a process including paper texture masking, ink bleeding, gamma and rescaling. Also, the clients are set to send sparse updates of the most significant gradient updates to the server. The total number clients do not directly affect the attack vector, see Process-II in Fig. 2, but indirectly through update scheduling.
@@ III-D RL Agent Design
We propose a PPO-based agent to generate adversarial waveforms or image perturbations that consistently induce parameter updates clustered within targeted indices of the target models’ parameter vector, denoted by $\bm{\theta}\in\mathbb{R}^{M}$. The agent learns under the guidance of client parameter updates.
@@ III-D1 Observation and Action Spaces
  • Observation: $\mathcal{O}=\{x_{\text{clean}},\mathbf{u}_{t}\}$, with binary vector $\mathbf{u}_{t}\in\{0,1\}^{M}$ indicating recently updated parameters.
  • Action: $\mathcal{A}=\{\delta\in\mathbb{R}^{D}\mid\|\delta\|_{\infty}\leq\epsilon\}$, constrained adversarial perturbations. To improve learning efficiency, action space of the agent is compressed with a patch encoder. The input perturbations are parameterized in a low-dimensional latent patch space. For ASR systems, the patches are a set of equal-length waveforms, and for the computer vision system, the patches are grids of sizes such as 8x8, 16x16. Then, a fixed decoder maps these patches to full-resolution perturbations, reducing dimensionality while preserving spatial structure. This latent representation serves as a unified action space for reinforcement learning.
@@ III-D2 Reward Formulation
Target Specificity: Ratio term emphasizes updates within targeted parameter indices $[i:j]$, here, the target is acquired not by designation, but rather condensed from observed preliminary dominant indices in early rounds.
Perceptibility Penalty: For ASR targets, the perceptual constraint simultaneously limits frequency-domain artifacts and temporal waveform distortions [10], calculated as follows.
$\mathcal{L}_{\text{perc}}=\lambda_{1}\|\text{STFT}(\mathcal{A}+x_{\text{clean}})-\text{STFT}(x_{\text{clean}})\|_{F}+\lambda_{2}\,\text{rms}(\mathcal{A})$ (8)
where STFT is the Short-Time Fourier Transform, rms is root mean square, and $\lambda_{1,2}$ are weight factors.
For CV targets, the constraint is $\mathcal{L}_{\text{perc}}=\lambda\|\mathcal{A}\|_{2}^{2}$.
@@ III-D3 Policy Network Architecture
The policy network consists of a target specific feature extractor and multi-layer perceptron decoders. The feature extractor shares the typical structures of the model of the target system. For the ASR targets, the extractor first processes raw waveforms through a multi-resolution convolutional frontend, followed by bidirectional temporal modelling, before diverging into stage-dependent perturbation generators, while for the computer vision target, the extractor consists of convolution blocks like that of a ResNet model [43].
@@ IV-A Experimental Setups
$\text{$BW$}=\frac{\text{Data Rate}\times\text{Bit-Width}}{8\text{ bits/byte}}$ (9)
According to equation (9), we can first get the bandwidth of tested module is 18.75 GB/s. Therefore, the data throughput for each 64ms refresh period will be approximately 1.2 GB. Furthermore, the default tRC is $46.16~\text{ns}$.
In an operating system-level attack environment, the optimal ACT rate based on the clflushopt instruction sequence is 159 ACTs/tREFI, meaning that the actual average tRC used for frequent activations is $49~\text{ns}$, which is close to the theoretical ACT limit. Under this condition, the attacker can issue up to 1,306K activations per refresh window. Moreover, all other timing parameters of tested module follow the default configuration.
@@ IV-B Rowhammer Test Bench
Considering that FL may use different memory replacement strategies and that data diversity plays a role, this work tested different Rowhammer attack thresholds, based on two access patterns and four data patterns. Compare to single-sided access pattern, the double-sided requires less activations. However, it is also more difficult to utilise as it requires two adjacent rows for victim data. The data pattern will be the data layout between victim rows and aggressor rows. The tests not only reveal patterns that are more likely to induce bit flips, such as 1 hammered by 0 or 1 hammered by 0, but also demonstrate relatively more challenging patterns, such as 0 hammered by 0 or 1 hammered by 1. Specially, we use hexadecimal strings to represent data patterns that fills the entire DRAM row. For example, 0xFF represents 32-bit data filled with all ones, while 0x00 represents all zeros. The obtained attack thresholds are derived from the average values across all banks of the tested DRAM module.
These activation thresholds ensure that the attack induces bit-flips in at least 95% of the DRAM rows. The inability to obtain thresholds for all rows is likely due to the presence of exceptionally resilient cells within the module, which results in extreme deviations in the average threshold. These thresholds are specifically presented in Table II. All threshold values fall within the theoretical limit of 1,309K activations and reserve enough dummy rows activations to bypass TRR. It can be observed that double-sided Rowhammer generally requires fewer activations than single-sided one. In the worst case, 265K activations are sufficient to induce bit flips.
@@ IV-C PPO Agent Training
TABLE II: Rowhammer activation thresholds under different data patterns.
Victim Aggressor Single-sided Double-sided
0xFF 0x00 185K 115K
0x00 0xFF 240K 140K
0x55 0x55 260K 160K
0xAA 0xAA 265K 165K
@@
We empirically evaluate the effectiveness of our proposed PPO-based adversarial framework in consistently inducing clustered and repetitive parameter updates in target FL systems under varying sparsity settings, which are suitable for Rowhammer exploitation.
IV-C1 FL Setup
@@
We selected three widely-used ASR models of varying complexity, parameter counts, and architectural patterns for evaluation and a MobileNetV3 model, see in Table I the four models and their parameter statistics.
@@
For all models, we simulate for two sparsity settings: 1) 0.1% updates of moderate communication efficiency is required without severely limiting model convergence speed; and 2) 0.05% updates that seek extreme communication efficiency conditions, simulating highly bandwidth-constrained or large-scale deployments.
The total number of simulated clients is training efficiency oriented, and scales subject to limited hardware capability and target model sizes, ranging from 20-50.
Also, to improve training efficiency, all target models are quantized version, as listed in I.
@@
The PPO agent was trained to generate adversarial waveforms constrained by a maximum perturbation limit ($\|\delta\|_{\infty}\leq 0.1$), ensuring the adversarial audio samples remain stealthy and practically imperceptible. For the ASR models, the training dataset employed was the English subset of Common Voice 17.0 [40], and for CV, the dataset is CIFAR10 [41]. Hyperparameters for the reward function were empirically started with $\alpha=1.0$, $\beta=0.8$, and $\gamma=0.6$, with perceptibility penalty weights set to $\lambda_{1}=\lambda_{2}=0.5$ for ASR, and $\lambda=0.8$ for CV.
IV-C2 Training Protocol
The target FL system, serving as the environment for the RL agent, at each iteration, is set to run 100 communication rounds before being reset at the end. The agent generates action and observe client updates at each round. The agent is train for a total of 800 such iterations.
TABLE III: Estimated Row Activations Based on Empirical Clustering
Model Sparsity $H_{\text{max}}$ RUR(%) $\mathbb{E}_{Act}$
Conformer-CTC-S 0.1% 296K 69.5 206K
@INT4 0.05% 592K 60.5 358K
Squeezeformer-XS 0.1% 286K 63.1 180K
@INT4 0.05% 572K 53.7 307K
QuartzNet 5x5 0.1% 192K 76.0 145K
@INT8 0.05% 384K 60.8 233K
MobileNetV3 0.1%  [^4] K 63.3 281K
@INT8 0.05%  [^8] K 58.3 518K
V Results and Analysis
V-A RL Performance
Table III and Fig. 4 present results of our PPO method across different model architectures and sparsity levels. On the target of Repeated Update Rate (RUR), as summarized in Table III, our PPO agent consistently uplifts the model update concentration at both sparsity settings. At high sparsity (0.05%), the agent achieved RUR ranging from $53.7\%$ (Squeezeformer-XS) to $60.8\%$ (QuartzNet 5x5); while at medium sparisty (0.1%), RUR are higher, ranging from $63.1\%$ (Squeezeformer-XS) to $76.0\%$ (QuartzNet 5x5). These results demonstrate the efficacy of our approach in inducing concentrated and stable update patterns necessary for facilitating targeted Rowhammer attacks in realistic FL settings.
Figure 4: Repeated Update Rate of target models with sparsity of 0.1% and 0.05% settings, in all cases, the agent succeeded in inducing repetitive clustered updates. At sparsity of 0.1%, attacks on all models achieved RUR above 60% with QuartzNet 5x5 reached 76%. At sparsity of 0.05%, the RUR of these attacks decreased, but remain above 50%, with the highest QuartzNet 5x5 attack at 60.8%.
V-B Bit-Flip Validation